# A Multimodal Data Fusion Generative Adversarial Network for Real Time Underwater Sound Speed Field Construction

### v1

@@ 2.3 Discriminator
Here we use the coordinates of the target point and the SST as the output of the discriminator, which is used as an alternative to the traditional way of training the discriminator in terms of realism. The input tensor is stacked with the same multi-layer stack of cross-modal attention blocks and position-wise feed-forward networks used in the generator, where the Query is the SSP feature vector of the target point $\mathbf{Z}_{ssp}^{t}$, the Key is the SSP feature vector of the reference point $\mathbf{Z}_{ssp}^{r,n}$ , and the Value corresponds to the feature vector encoded by the coordinates of the reference point, and the feature vector encoded by the SST $\mathbf{Z}_{loc}^{r,n}$ and $\mathbf{Z}_{sst}^{r,n}$ , respectively. The formula for each layer of attention is similar to the computational strategy of the generator:
Figure 3: Visualization of intermediate features in the MDF-RAGAN model. (a)-(c) represent the intermediate outputs from the generator, while (d)-(f) show the intermediate feature outputs from the discriminator. Specifically: (a) represents the fused label features of the target point, (b) represents the SSP features of surrounding sample points, (c) represents the fused label features of surrounding sample points, (d) represents the fused label features of the target point, (e) represents the fused label features of surrounding sample points, and (f) represents the SSP features of surrounding sample points.
$O^{\left(l\right)}=\text{Soft}\max\left(\frac{Q^{\left(l-1\right)}K^{\left(l-1\right)^{\mathbf{T}}}-m^{\left(l\right)}}{\sqrt{d}}\right)V^{\left(l-1\right)},$ (18)
[…]
@@
$\mathbf{\hat{t}}_{sst}=W_{sst}H_{sst}^{\left(l\right)}+b_{sst},$ (21)
The above multi-task outputs, without adding extra parameters, enable the discriminator to both improve its sensitivity to differences in SSP details and provide the generator with error signals for both coordinates and SST, promoting stable convergence and improving the accuracy of the overall network in complex underwater acoustic environments. Fig. 3 visualizes the fused feature of generator and discriminator.
2.4 Optimization Strategy
@@
In order to guarantee the stability and convergence effect of MDF-RAGAN in the early stage of adversarial training, we adopt a staged training strategy, in the first stage, we only optimise the generator parameters $\theta_{G}$ , the discriminator parameters keep $\theta_{D}$ kept frozen and not updated, at this time through the
$\theta_{G}\leftarrow\theta_{G}-\eta_{t}\,\nabla_{\theta_{G}}\mathcal{L}_{G}.$ (31)
In the first stage, the generator learns an initial mapping from the multi-source fused label embeddings to the SSP perturbations at the target point. A linear learning-rate warm-up is applied, and the network parameters are updated solely with the generator loss $\mathcal{L}_{G}$. In the second stage, training switches to adversarial learning: network parameters are alternately updated with the generator loss $\mathcal{L}_{G}$ and the discriminator loss $\mathcal{L}_{D}$, enabling joint optimization of both networks.
@@
3 Results and Discussions
3.1 Experimental Setting
3.1.1 Dataset
@@
SSP data were taken from the GDCSM_Argo gridded Argo dataset (Li Hong, 2017), and the raw observation profiles covered 0-1976 m depth with 58 scattered points, which were linearly interpolated depth-by-depth to 1 m intervals to generate 1977-dimensional high-resolution SSP vectors. SST data were obtained from the NOAA OISST monthly average product (Huang et al., 2021a), spatially mapped and aligned to an Argo grid to obtain the current month’s SST values for each grid point.
@@
3.1.2 Performance Evaluation Metrics
@@ 3.1.3 Parameter Initialization
3.2 Baselines
In order to comprehensively validate the performance enhancement of the proposed MDF-RAGAN model in SSP estimation, three types of baseline methods are selected in this study: averaged profile based on historical statistics, spatial interpolation (SITP), and convolutional neural network (CNN) (Wu et al., 2024). Quantitative results show that MDF-RAGAN, by effectively fusing physical priors and modelling nonlinear interactions, consistently outperforms these baselines.
3.2.1 Mean Profile
@@ 3.2.2 Spatial Interpolation
where $D_{n}=\sqrt{\left(x-x_{n}\right)^{2}+\left(y-y_{n}\right)^{2}}$ , is the in-plane distance between the target point and the location of the first $n$ known profile, usually taken as a power exponent $p=2$ The SITP method weights historical observations within a geographic neighborhood, and can capture the effects of thermohaline gradients at spatially adjacent points to some extent.
3.2.3 CNN Model
As a deep learning baseline, this study implements a classical CNN model for comparing the performance of deep networks in SSP tasks (Wu et al., 2024). The model takes the center and eight profile vectors, corresponding coordinates and SST of its 3×3 neighborhood grid as multi-channel inputs, and extracts local features in the depth dimension through three-layer one-dimensional convolution (convolution kernel width of 3, channel numbers of 64, 128 and 256 in turn), ReLU activation and maximum pooling; subsequently, the perturbation value of the center is outputted by the regression of the two-layer fully-connected network. During the training process, the CNN is also optimized using the RMSE as the loss function.
3.3 Accuracy Performance
3.3.1 Influence of Training Data Length
@@
(a) Location 1: ($46.5^{\circ}$, $19.5^{\circ}$), 1975m
(b) Location 2: ($40.5^{\circ}$, $31.5^{\circ}$), 1975m
(c) Location 3: ($55.5^{\circ}$, $1.5^{\circ}$), 1975m
(d) Location 4: ($55.5^{\circ}$, $37.5^{\circ}$), 1975m
In addition, Table 1 compares the RMSE of MDF-RAGAN trained on 5, 10 and 17 years of historical data with the corresponding CNN, as well as SITP with the historical mean (MEAN) method, at four different depths (200 m, 400 m, 1000 m, 1975 m) and four typical shallow and deep sea coordinate points. Overall, the RMSE of MDF-RAGAN shows a continuous decrease at all depths and locations as the length of the training data expands from 5 to 17 years: for example, the RMSE of MDF-RAGAN at a depth of 200 m (point 1: ($58.5^{\circ}$, $19.5^{\circ}$)) decreases from 0.332 m/s to 0.236 m/s; at the same location at a depth of 1000 m, its error is more than halved from 0.187 m/s to 0.095 m/s. This shows that the longer historical window allows the model to learn a more adequate pattern of thermohaline change, especially in the deep ocean region, where the error decreases more significantly. Further comparing the different depths, it can be seen that the initial RMSE of MDF-RAGAN is higher than that of the deep sea (1000 m, 1975 m) in the shallow sea region (200 m and 400 m), but with the increase of the training years, its error rapidly approaches the deep sea level, and stays in the low range of 0.10-0.20 m/s in all depths at the 17-year window, showing a weak sensitivity to the depth, while the CNN and traditional methods, although performing better in the deep sea, consistently maintain a high amplitude error interval of 0.20-0.95 m/s at shallow depths. In following experiments, the training data length will be 17 years.
Table 1: Comparison of RMSE with different training data lengths
@@
MDF-RAGAN ($58.5^{\circ}$, $19.5^{\circ}$) 0.187 0.138 0.125 0.140 0.105 0.095
($58.5^{\circ}$, $26.5^{\circ}$) 0.168 0.151 0.128 0.128 0.114 0.100
($55.5^{\circ}$, $22.5^{\circ}$) 0.226 0.176 0.176 0.168 0.133 0.131
($52.5^{\circ}$, $22.5^{\circ}$) 0.266 0.242 0.245 0.199 0.181 0.182
CNN ($58.5^{\circ}$, $19.5^{\circ}$) 0.455 0.460 0.615 0.336 0.341 0.457
($58.5^{\circ}$, $26.5^{\circ}$) 0.545 0.445 0.572 0.408 0.326 0.427
($55.5^{\circ}$, $22.5^{\circ}$) 0.541 0.482 0.661 0.412 0.366 0.501
($52.5^{\circ}$, $22.5^{\circ}$) 0.648 0.542 0.613 0.487 0.410 0.465
SITP ($58.5^{\circ}$, $19.5^{\circ}$) 0.437 0.437 0.437 0.319 0.319 0.319
($58.5^{\circ}$, $26.5^{\circ}$) 0.194 0.194 0.194 0.151 0.151 0.151
($55.5^{\circ}$, $22.5^{\circ}$) 0.383 0.383 0.383 0.315 0.315 0.315
($52.5^{\circ}$, $22.5^{\circ}$) 0.385 0.385 0.385 0.357 0.357 0.357
MEAN ($58.5^{\circ}$, $19.5^{\circ}$) 0.427 0.427 0.427 0.318 0.318 0.318
($58.5^{\circ}$, $26.5^{\circ}$) 0.219 0.219 0.219 0.164 0.164 0.164
($55.5^{\circ}$, $22.5^{\circ}$) 0.457 0.457 0.457 0.373 0.373 0.373
($52.5^{\circ}$, $22.5^{\circ}$) 0.503 0.503 0.503 0.484 0.484 0.484
@@
Table 2: RMSE comparison of different methods at 1976 m depth across 10 typical locations (data from 2021.01-2023.06)
Number Location MDF-RAGAN (m/s) CNN (m/s) MEAN (m/s) SITP (m/s)
1 ($58.5^{\circ}$, $1.5^{\circ}$) 0.0673 0.3443 0.3120 0.2211
2 ($49.5^{\circ}$, $13.5^{\circ}$) 0.1191 0.2700 0.3028 0.2050
3 ($49.5^{\circ}$, $16.5^{\circ}$) 0.1339 0.2902 0.2921 0.1991
4 ($49.5^{\circ}$, $19.5^{\circ}$) 0.2151 0.3527 0.2891 0.2123
5 ($49.5^{\circ}$, $22.5^{\circ}$) 0.1877 0.3710 0.3458 0.2883
6 ($49.5^{\circ}$, $31.5^{\circ}$) 0.0948 0.2203 0.4415 0.3754
7 ($46.5^{\circ}$, $10.5^{\circ}$) 0.1082 0.2157 0.5471 0.5558
8 ($46.5^{\circ}$, $37.5^{\circ}$) 0.1543 0.2287 0.2581 0.2016
9 ($43.5^{\circ}$, $22.5^{\circ}$) 0.2180 0.2762 0.3906 0.3745
10 ($43.5^{\circ}$, $37.5^{\circ}$) 0.1730 0.3237 0.7380 0.5080
Average 0.1472 0.2893 0.3917 0.3141
3.3.2 Accuracy in Deep and Shallow Water
To evaluate the accuracy performance in estimating the SSP, we compare the results of MDF-RAGAN with mean profile, SITP and CNN on 10 typical latitude and longitude coordinate points (30 months of data from 2021.01-2023.06) as shown in Table 2. The results show that the RMSE of the historical average MEAN method fluctuates around 0.25$\sim$0.73 m/s, and the RMSE of SITP fluctuates around 0.19$\sim$0.55 m/s. While CNN compresses the results to around 0.21$\sim$0.37 m/s, the MDF-RAGAN proposed in this paper focuses on 0.06$\sim$0.21 m/s, and the mean value is only 0.147 m/s, which is 62% higher than the traditional mean method, and more than 53% compared with the spatial interpolation method. Even compared with CNN, the performance is improved by 49%. The table fully demonstrates the accuracy and stability of MDF-RAGAN for estimation in a large-scale sound velocity gradient environment.
(a) ($49.5^{\circ}$, $13.5^{\circ}$), 200m
(b) ($49.5^{\circ}$, $28.5^{\circ}$), 200m
(c) ($52.5^{\circ}$, $25.5^{\circ}$), 200m
(d) ($49.5^{\circ}$, $13.5^{\circ}$), 400m
(e) ($49.5^{\circ}$, $28.5^{\circ}$), 400m
(f) ($52.5^{\circ}$, $25.5^{\circ}$), 400m
(g) ($49.5^{\circ}$, $13.5^{\circ}$), 1000m
(h) ($49.5^{\circ}$, $28.5^{\circ}$), 1000m
(i) ($52.5^{\circ}$, $25.5^{\circ}$), 1000m
(j) ($49.5^{\circ}$, $13.5^{\circ}$), 1975m
(k) ($49.5^{\circ}$, $28.5^{\circ}$), 1975m
(l) ($52.5^{\circ}$, $25.5^{\circ}$), 1975m

### v3

@@ II-C Discriminator
Here we use the coordinates of the target point and the SST as the output of the discriminator, which is used as an alternative to the traditional way of training the discriminator in terms of realism. The input tensor is stacked with the same multi-layer stack of cross-modal attention blocks and position-wise feed-forward networks used in the generator, where the Query is the SSP feature vector of the target point $\mathbf{Z}_{ssp}^{t}$, the Key is the SSP feature vector of the reference point $\mathbf{Z}_{ssp}^{r,n}$, and the Value consists of the feature vectors encoded by the reference point coordinates $\mathbf{Z}_{loc}^{r,n}$ and the SST $\mathbf{Z}_{sst}^{r,n}$, respectively. The discriminator applies this attention mechanism with identical numerical stabilization as described in Section II-B1 to enforce physical consistency between the generated SSP and the associated environmental conditions.
@@
$\mathbf{\hat{t}}_{sst}=W_{sst}H_{sst}^{\left(l\right)}+b_{sst}.$ (15)
In contrast to conventional GAN discriminators that output a single real/fake logit, the discriminator in MDF-RAGAN is formulated as a multi-task regressor that explicitly predicts the SST and geographic coordinates associated with each candidate SSP. This design yields richer and smoother gradient signals than a one-bit decision, since the regression logits encode physically meaningful environmental constraints tied to the underlying oceanographic state. By forcing the discriminator to recover these variables, the adversarial learning process enforces cross-modal physical consistency and suppresses solutions that are acoustically plausible but environmentally unrealistic. As a result, the discriminator serves not only as a real/fake judge but also as a validator of oceanographic coherence, which stabilises training and enhances the fidelity of the reconstructed sound-speed perturbations in complex marine environments.
II-D Optimization Strategy
@@
In order to guarantee the stability and convergence effect of MDF-RAGAN in the early stage of adversarial training, we adopt a staged training strategy. In the first stage, we only optimise the generator parameters $\theta_{G}$, while the discriminator parameters $\theta_{D}$ are kept frozen and not updated. The generator learns an initial mapping from the multi-source fused label embeddings to the SSP perturbations at the target point. A linear learning-rate warm-up is applied, and the network parameters are updated solely with the generator loss $\mathcal{L}_{G}$. In the second stage, training switches to adversarial learning: network parameters are alternately updated with the generator loss $\mathcal{L}_{G}$ and the discriminator loss $\mathcal{L}_{D}$, enabling joint optimization of both networks.
@@
III Results and Discussions
III-A Experimental Setting
III-A1 Dataset
@@
SSP data were taken from the GDCSM_Argo gridded Argo dataset [26], and the raw observation profiles covered 0-1976 m depth with 58 scattered points, which were linearly interpolated depth-by-depth to 1 m intervals to generate 1977-dimensional high-resolution SSP vectors. SST data were obtained from the NOAA OISST monthly average product [27], spatially mapped and aligned to an Argo grid to obtain the current month’s SST values for each grid point.
@@
III-A2 Performance Evaluation Metrics
@@ III-A3 Parameter Initialization
III-B Baselines
In order to comprehensively validate the performance enhancement of the proposed MDF-RAGAN model in SSP estimation, three types of baseline methods are selected in this study: averaged profile based on historical statistics, spatial interpolation (SITP), and convolutional neural network (CNN) [24]. Quantitative results show that MDF-RAGAN, by effectively fusing physical priors and modelling nonlinear interactions, consistently outperforms these baselines.
III-B1 Mean Profile
@@ III-B2 Spatial Interpolation
where $D_{n}=\sqrt{\left(x-x_{n}\right)^{2}+\left(y-y_{n}\right)^{2}}$ , is the in-plane distance between the target point and the location of the first $n$ known profile, usually taken as a power exponent $p=2$ . The SITP method weights historical observations within a geographic neighborhood, and can capture the effects of thermohaline gradients at spatially adjacent points to some extent.
III-B3 Attention-Assisted CNN Model
As a deep learning baseline, this study implements Att-CNN model for comparing the performance of deep networks in SSP tasks. The Att-CNN model is an improved version of the CNN architecture proposed in [24], incorporating attention mechanisms to capture long-range dependencies in the depth dimension [28]. Compared with the SOM-based method [29] which reconstructs SSP from remote sensing data using self-organizing maps, the Att-CNN demonstrates superior performance and represents the current state-of-the-art approach for SSP estimation. The model takes the center and eight profile vectors, corresponding coordinates and SST of its 3×3 neighborhood grid as multi-channel inputs, and extracts local features in the depth dimension through three-layer one-dimensional convolution (convolution kernel width of 3, channel numbers of 64, 128 and 256 in turn), ReLU activation and maximum pooling; subsequently, the perturbation value of the center is outputted by the regression of the two-layer fully-connected network. During the training process, the CNN is also optimized using the RMSE as the loss function.
III-C Accuracy Performance
III-C1 Influence of Training Data Length
@@
TABLE I: Comparison of RMSE with different training data lengths
@@
MDF-RAGAN ($58.5^{\circ}$S, $19.5^{\circ}$E) 0.187 0.138 0.125 0.140 0.105 0.095
($58.5^{\circ}$S, $26.5^{\circ}$E) 0.168 0.151 0.128 0.128 0.114 0.100
($55.5^{\circ}$S, $22.5^{\circ}$E) 0.226 0.176 0.176 0.168 0.133 0.131
($52.5^{\circ}$S, $22.5^{\circ}$E) 0.266 0.242 0.245 0.199 0.181 0.182
Att-CNN ($58.5^{\circ}$S, $19.5^{\circ}$E) 0.455 0.460 0.615 0.336 0.341 0.457
($58.5^{\circ}$S, $26.5^{\circ}$E) 0.545 0.445 0.572 0.408 0.326 0.427
($55.5^{\circ}$S, $22.5^{\circ}$E) 0.541 0.482 0.661 0.412 0.366 0.501
($52.5^{\circ}$S, $22.5^{\circ}$E) 0.648 0.542 0.613 0.487 0.410 0.465
SITP ($58.5^{\circ}$S, $19.5^{\circ}$E) 0.437 0.437 0.437 0.319 0.319 0.319
($58.5^{\circ}$S, $26.5^{\circ}$E) 0.194 0.194 0.194 0.151 0.151 0.151
($55.5^{\circ}$S, $22.5^{\circ}$E) 0.383 0.383 0.383 0.315 0.315 0.315
($52.5^{\circ}$S, $22.5^{\circ}$E) 0.385 0.385 0.385 0.357 0.357 0.357
MEAN ($58.5^{\circ}$S, $19.5^{\circ}$E) 0.427 0.427 0.427 0.318 0.318 0.318
($58.5^{\circ}$S, $26.5^{\circ}$E) 0.219 0.219 0.219 0.164 0.164 0.164
($55.5^{\circ}$S, $22.5^{\circ}$E) 0.457 0.457 0.457 0.373 0.373 0.373
($52.5^{\circ}$S, $22.5^{\circ}$E) 0.503 0.503 0.503 0.484 0.484 0.484
(a) Location 1: ($55.5^{\circ}$S, $1.5^{\circ}$E), 1975m
(b) Location 2: ($55.5^{\circ}$S, $37.5^{\circ}$E), 1975m
In addition, Table I compares the RMSE of MDF-RAGAN trained on 5, 10 and 17 years of historical data with the corresponding CNN, as well as SITP with the historical mean (MEAN) method, at four different depths (200 m, 400 m, 1000 m, 1975 m) and four typical shallow and deep sea coordinate points. Overall, the RMSE of MDF-RAGAN shows a continuous decrease at all depths and locations as the length of the training data expands from 5 to 17 years: for example, the RMSE of MDF-RAGAN at a depth of 200 m (point 1: ($58.5^{\circ}$S, $19.5^{\circ}$E)) decreases from 0.332 m/s to 0.236 m/s; at the same location at a depth of 1000 m, its error is more than halved from 0.187 m/s to 0.095 m/s. This shows that the longer historical window allows the model to learn a more adequate pattern of thermohaline change, especially in the deep ocean region, where the error decreases more significantly. Further comparing the different depths, it can be seen that the initial RMSE of MDF-RAGAN is higher than that of the deep sea (1000 m, 1975 m) in the shallow sea region (200 m and 400 m), but with the increase of the training years, its error rapidly approaches the deep sea level, and stays in the low range of 0.10-0.20 m/s in all depths at the 17-year window, showing a weak sensitivity to the depth, while the CNN and traditional methods, although performing better in the deep sea, consistently maintain a high amplitude error interval of 0.20-0.95 m/s at shallow depths. In following experiments, the training data length will be 17 years.
@@
TABLE II: RMSE comparison of different methods at 1976 m depth across 10 typical locations (data from 2021.01-2023.06) (m/s)
Location MDF-RAGAN Att-CNN MEAN SITP
($58.5^{\circ}$S, $1.5^{\circ}$E) 0.067 0.344 0.312 0.221
($49.5^{\circ}$S, $13.5^{\circ}$E) 0.119 0.270 0.303 0.205
($49.5^{\circ}$S, $16.5^{\circ}$E) 0.134 0.290 0.292 0.199
($49.5^{\circ}$S, $19.5^{\circ}$E) 0.215 0.353 0.289 0.212
($49.5^{\circ}$S, $22.5^{\circ}$E) 0.188 0.371 0.346 0.288
($49.5^{\circ}$S, $31.5^{\circ}$E) 0.095 0.220 0.442 0.375
($46.5^{\circ}$S, $10.5^{\circ}$E) 0.108 0.216 0.547 0.556
($46.5^{\circ}$S, $37.5^{\circ}$E) 0.154 0.229 0.258 0.202
($43.5^{\circ}$S, $22.5^{\circ}$E) 0.218 0.276 0.391 0.375
($43.5^{\circ}$S, $37.5^{\circ}$E) 0.173 0.324 0.738 0.508
Average 0.147 0.289 0.392 0.314
III-C2 Influence of Coverage Density
To evaluate the sensitivity of SSP reconstruction to the spatial coverage density of historical profiles, we enlarge the original $3\times 3$ neighbourhood to a sparser $5\times 5$ window while still selecting nine reference profiles. This configuration increases the average spacing between reference points, weakens local spatial correlations, and thus provides a more challenging sparse-coverage scenario. As summarised in Table III, all methods exhibit increased RMSEs under this setting, indicating that denser historical sampling is beneficial for accurate SSP reconstruction.
The degree of performance degradation, however, varies substantially across models. MEAN and SITP show clear error inflation at all depths, and the Att-CNN baseline also suffers noticeable accuracy loss, especially within the upper 400 m where thermocline variability is strongest. By contrast, MDF-RAGAN consistently achieves the lowest RMSE across all depth layers and locations: even with sparse neighbourhoods, its integrated-profile errors over 0–1000 m and 0–1975 m remain at 0.459 and 0.336 m/s, compared with 0.653/0.482 m/s for Att-CNN and 0.964/0.745 m/s for SITP (Table III), and it maintains a clear advantage at representative stations along the meridional transect (Table IV).
TABLE III: RMSE (m/s) of different methods at various depths under the sparse-coverage setting, characterising the effect of vertical coverage density of historical SSPs.
Depth (m) MDF-RAGAN Att-CNN MEAN SITP
200 0.632 0.952 1.591 1.391
400 0.575 0.844 1.458 1.241
1000 0.459 0.653 1.168 0.964
1975 0.336 0.482 0.907 0.745
To further examine the impact of spatial coverage density, we analyse a meridional transect where historical sampling gradually decreases from high to lower latitudes, spanning $(57.5^{\circ}\mathrm{S},\,22.5^{\circ}\mathrm{E})$ to $(42.5^{\circ}\mathrm{S},\,2.5^{\circ}\mathrm{E})$. Table IV summarises the RMSEs of each method at 200 m, 400 m, and 1000 m for four representative locations. As coverage becomes sparser equatorward, all methods show error growth, but MDF-RAGAN degrades more gently than Att-CNN, MEAN, and SITP, and keeps RMSE values below 0.8 m/s at all depths and stations.
TABLE IV: RMSE (m/s) of different methods at representative locations with varying historical coverage density.
($57.5^{\circ}$S, $22.5^{\circ}$E) ($52.5^{\circ}$S, $22.5^{\circ}$E) ($47.5^{\circ}$S, $22.5^{\circ}$E) ($42.5^{\circ}$S, $2.5^{\circ}$E)
Method 200 m 400 m 1000 m 200 m 400 m 1000 m 200 m 400 m 1000 m 200 m 400 m 1000 m
MDF-RAGAN 0.477 0.390 0.274 0.702 0.631 0.521 0.720 0.674 0.521 0.556 0.535 0.412
Att-CNN 1.001 0.801 0.561 0.916 0.803 0.622 1.052 0.934 0.740 0.863 0.751 0.562
MEAN 3.229 2.553 1.945 2.087 1.548 1.082 1.114 1.259 1.015 0.768 1.125 0.881
SITP 2.819 2.256 1.788 1.493 1.106 0.790 0.857 0.892 0.681 0.900 0.890 0.659
The robustness of MDF-RAGAN under sparse coverage mainly stems from two aspects. First, the residual perturbation formulation allows the generator to preserve the large-scale background structure while only correcting fine-scale deviations, even when reference profiles are widely spaced. Second, the cross-modal perturbation attention aggregates long-range, SST and location-consistent information from historical profiles, suppressing the impact of locally missing observations. As a result, MDF-RAGAN exhibits the smallest performance degradation among all compared methods and remains the most accurate approach under sparse-sampling conditions, which is crucial for large-scale ocean acoustic applications in data-poor regions.
III-C3 Accuracy in Deep and Shallow Water
[…]
(a) ($49.5^{\circ}$S, $13.5^{\circ}$E), 200m
(b) ($49.5^{\circ}$S, $28.5^{\circ}$E), 200m
(c) ($52.5^{\circ}$S, $25.5^{\circ}$E), 200m
(d) ($49.5^{\circ}$S, $13.5^{\circ}$E), 400m
(e) ($49.5^{\circ}$S, $28.5^{\circ}$E), 400m
(f) ($52.5^{\circ}$S, $25.5^{\circ}$E), 400m
(g) ($49.5^{\circ}$S, $13.5^{\circ}$E), 1000m
(h) ($49.5^{\circ}$S, $28.5^{\circ}$E), 1000m
(i) ($52.5^{\circ}$S, $25.5^{\circ}$E), 1000m
(j) ($49.5^{\circ}$S, $13.5^{\circ}$E), 1975m
(k) ($49.5^{\circ}$S, $28.5^{\circ}$E), 1975m
(l) ($52.5^{\circ}$S, $25.5^{\circ}$E), 1975m